OpenStudy (anonymous):

solve sec x csc x - 2 csc x = 0

OpenStudy (turingtest):

factor out csc x and solve like you would a normal polynomial

OpenStudy (anonymous):

When you factor out the cscx did you get secx cscx cot^2x?

OpenStudy (turingtest):

\[\sec x\csc x-2\csc x=0\]\[\csc x(\sec x-2)=0\]now you just have to solve\[\csc x=0\text{ and }\sec x-2=0\]

OpenStudy (anonymous):

Thanks, I can get it from here!